Operations Report – April 17th

Crew 210 Operations Report 17-APR-2019

SOL: 5

Name of person filing report: Shannon Rupert

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run last night): 10 (7:30 am to 9:30 pm)

Solar— SOC 79 % when generator was turned on last night

Diesel Reading – 38 %

Propane Reading – 49 % (Note: I believe I misread this earlier)

Ethanol Free Gasoline – 4 gallons.

Water (auxillary tank) – 0 gallons

Water (static tank) – 300 gallons

Auxillary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – unknown gallons.

Water (loft) – Static to Loft Pump used – no

Water Meter: not read

Toilet tank emptied: no

Deimos rover used: Still non-operational


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, yes

Hours: Did not record

Beginning charge: 100

Ending charge: 96

Currently charging: yes

Spirit rover used: yes

Hours: did not check

Beginning charge: 100

Ending charge: 90

Currently charging: yes

Opportunity rover used: yes

Hours: not recorded

Beginning charge: 100

Ending charge: 94

Currently charging: yes

Curiosity rover used: No

Hours: 88.8

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) Honda, 300, 350.1, 350.3

Reason for use: Testing and Field work

Oil Added? No

ATV Fuel Used: 3 Gals.

# Hours the ATVs were Used today:4

Notes on ATVs: Battery in the Honda appears to be going downhill. We
will keep an eye on it and take jumper cables when we use it.

HabCar used and why, where? No

CrewCar used and why, where? Yes, to Hanksville for curry and a test
for live online event tomorrow

General notes and comments: We need a new hose for the compressor.
Current one has several leaks that have been taped.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: David will send a full report today.

Summary of ScienceDome operations: Things are rockin’ in the lab.
Plant work and soil work in progress.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: First aid kits have been used and
so not all have the necessary supplies to be useful. They should be
gone through and restocked. I will try to get this done before the
next crew arrives.

Questions, concerns and requests to Mission Support: Nothing tonight,
thank you.

Copyright © The Mars Society. All rights reserved. | Main Site